# Fairness in Social Influence Maximization via Optimal Transport

### v3

@@ 1 Introduction
Social networks play a fundamental role in the spread of information, as in the context of commercial products endorsement [18], job vacancy advertisements [3], public health awareness [27], etc. Information, ideas, or new products can either go viral and potentially bring significant changes in a community or die out quickly. In this context, a fundamental algorithmic problem arises, known as Social Influence Maximization (SIM) [12, 13]. SIM studies how to strategically select a pre-specified small proportion of nodes in the social network, the early adopters or seeds so that the outreach generated by a diffusion process that starts at these early adopters is maximized. Consider, for example, a product endorsement campaign: the early adopters are strategically selected users who receive the product first to promote it to their friends, who in turn may or may not adopt it. The optimal selection of early adopters is known to be an NP-hard problem [12]. Thus, many heuristic strategies have been proposed, based on iterative processes such as greedy algorithms or on network centrality measures. However, all these algorithms purely rely on the graph topology and are agnostic to users’ demographics, which raises significant fairness concerns, especially in contexts of health awareness campaigns, education, and job advertisements, where one wants to ensure an equitable spreading of information. Indeed, real-world social networks are populated by different social groups, based on gender, age, race, geography, etc., with different group sizes or connectivity patterns. Ignoring these aspects and focusing only on the outreach maximization process usually leads to the early adopters being the most central nodes. Consequently, low-interconnected minorities are often neglected from the diffusion process, thus causing fundamental inequity in the information propagation and biases exacerbation [11, 25].
@@
The problem of SIM was first introduced in 2003 in Kempe et al. [12], where the problem of optimally selecting a (limited) set of early adopters was proved to be NP-hard. The study of SIM under fairness guarantees has a more recent history [6]. Several multiple group-level metrics of fairness have been proposed over the years [7]. They fall under the notions of equity [23, 10, 11], equality [7], max-min fairness [8, 30], welfare [17], and diversity [25]: all of them quantify the fair distribution of influence across groups. In particular, Stoica et al. [23] propose a new SIM algorithm that operates under the constraint that, in expectation, the same percentage of users in each category is reached. Junaid et al. [10] optimize outreach under fairness and time constraints, by ensuring that the expected fraction of influenced nodes in each group is the same within a prescribed time deadline. Farnadi et al. [7] propose a unifying framework that encodes all different definitions of fairness in the SIM process as constraints in a linear program that optimizes outreach. Several other works [8, 30] adopt a max-min strategy. Specifically, in Fish et al. [8] fairness is ensured by maximizing the minimum probability of a group receiving the information through modifications of the greedy algorithm. Zhu et al. [30] ensure that the outreach contains a pre-specified proportion of each group in a population. Finally, Tsang et al. [25] optimize outreach under the constraint that no group should be better off by leaving the influence maximization process with their proportional allocation of resources done internally. All these definitions involve a marginal expected value of fairness in groups, without considering the correlations – or other higher-order moments – for the joint probability distribution of different groups adopting the information (see Farnadi et al. [7] for an overview). In contrast, our work introduces a novel formalism for taking into account the actual joint distribution of outreach among groups, thus considering all groups simultaneously, highlighting limitations of various fairness metrics and developing a new seed selection policy that strategically extracts and optimizes our proposed notion of fairness. Finally, our work is inspired by a recent line of work that draws on optimal transport theory [28] for fairness guarantees [2, 4, 21, 29, 20, 24]. To our knowledge, this is the first work to develop novel metrics and seeding algorithms that leverage optimal transport for the SIM problem.
@@
Many models of diffusion processes in the SIM problem are inherently stochastic, meaning that who gets the information transmitted can vary greatly from one run to another. Consider, as an example, the case in which $50\%$ of realizations over a diffusion process, no one in group 1 receives the information and everyone in group 2 does, whereas in the other $50\%$ it is the opposite. This circumstance would be classified as fair in expectation, even though it is commonly not perceived as “fair”. We show how this phenomenon is also common in real-world data and how our proposed framework can detect such undesired scenarios. This prompts us to study a novel fairness metric.
@@ 2 Preliminaries
For a network $G$ with two communities $(C_{i})_{i\in~{}[2]}$ and a seedset $S$, we let $x_{i}$, $i~{}\in~{}[2]$, denote the fraction of nodes in each community in the outreach $\phi_{G}(S)$. The final configuration is the tuple $(x_{1},x_{2})$.
In many definitions in the literature, fairness is operationalized as measuring the expected value of the final configuration, where the expectation is taken over the diffusion process. In particular, the equity definition introduced by Stoica et al. [23], Junaid et al. [10] checks that the expected value of the proportions of each group reached in the outreach is the same for all groups. For a formal definition of equity and other fairness definitions in the literature, see Appendix A. We will show that relying solely on the expected value leads largely unfair outcomes to be classified as fair.
@@ 3 Mutual Fairness via Optimal Transport
Consider the SIM problem with nodes belonging to two groups, $C_{1}$ and $C_{2}$, each group having the outreach probability distribution $\mu_{i}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{1},i\in\{1,2\}$, with $\delta_{k}$ representing the delta distribution centered at $k\in[0,1]$. That is, in $50\%$ of the cases all members in group $i$ receive the information (i.e., we get $x_{i}=1.0$) and in $50\%$ of the cases no one in group $i$ receives the information (i.e., we get $x_{i}=0.0$). It is therefore tempting to say that this setting is fair since $\mu_{1}$ and $\mu_{2}$ coincide and therefore share the same expected value. We argue that this information does not suffice to claim fairness. Indeed, consider the two following probability distributions over the final configurations:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}=0.5\cdot\delta_{(0,0)}+0.5\cdot\delta_{(1,1)},\qquad{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}=0.25\cdot\delta_{(0,0)}+0.25\cdot\delta_{(1,1)}+0.25\cdot\delta_{(0,1)}+0.25\cdot\delta_{(1,0)}.$
with $\delta_{(i,j)}$, representing the delta distribution centered at $(i,j)\in[0,1]^{2}$. Interestingly, both ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ are “compatible” with $\mu_{1}$ and $\mu_{2}$: If we compute their marginals, we obtain $\mu_{1}$ and $\mu_{2}$. However, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ encode two fundamentally different final configurations. In ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$, the percentage of members of group 1 who get the information always coincides with the percentage of people of group 2. Conversely, in ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$, more outcomes are possible; in particular, there is a probability of $0.25+0.25=0.5$ that all members of one group receive the information and no member of the other group receives it (see Fig. 2). Thus, from a fairness perspective, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ encode very different outcomes. We therefore argue that a fairness metric should be expressed in terms of joint probability distribution $\gamma$, and not solely based on its marginals $\mu_{1}$ and $\mu_{2}$, as commonly done in the literature [23, 10].
@@ 3.1 A Fairness Metric Based on Optimal Transport
$W_{c}(\gamma_{a},\gamma_{b})=\min_{\pi\in\Pi(\gamma_{a},\gamma_{b})}\mathbb{E}_{(x_{1},x_{2}),(y_{1},y_{2})\sim\pi},[c((x_{1},x_{2}),(y_{1},y_{2}))]$ (1)
where $\Pi(\gamma_{a},\gamma_{b})$ is the set of probability distributions over $([0,1]\times[0,1])\times([0,1]\times[0,1])$ so that its first marginal is $\gamma_{a}$ and its second marginal is $\gamma_{b}$. Intuitively, the optimal transport problem quantifies the minimum transportation cost to morph $\gamma_{a}$ into $\gamma_{b}$ when transporting a unit of mass from $(x_{1},x_{2})$ to $(y_{1},y_{2})$ costs $c((x_{1},x_{2}),(y_{1},y_{2}))$. The optimization variable $\pi$ is called transportation plan and $\pi((x_{1},x_{2}),(y_{1},y_{2}))$ indicates the amount of mass at $(x_{1},x_{2})$ displaced to $(y_{1},y_{2})$. Thus, its first marginal has to be $\gamma_{a}(x_{1},x_{2})$ (that is, $(x_{1},x_{2})$ has to be transported to some $(y_{1},y_{2})$) and its second marginal must be $\gamma_{b}(y_{1},y_{2})$ (that is, the mass at $(y_{1},y_{2})$ has to arrive from some $(x_{1},x_{2})$). If the transportation cost $c$ is chosen to be a $p\geq 1$ power of a distance $d$, then $(W_{d^{p}}(\cdot,\cdot))^{1/p}$ is a distance on the space of probability distributions. When the probability distributions are discrete (or the space $[0,1]$ is discretized), the transportation problem (1) is a finite-dimensional linear program and can therefore be solved efficiently [16].
@@
To operationalize the optimal transport problem (1), we therefore need to define (i) a transportation cost and (ii) a reference distribution $\gamma^{\ast}$. To define the transportation cost, we start with the following two considerations. First, moving mass along the diagonal should have a cost of 00, as it does not affect fairness but only the efficiency (the proportion of population reached in respective groups). Second, moving mass orthogonally towards the diagonal should come at a price, since the difference in group proportion outreach between groups 1 and 2 decreases. We quantify this price as the Euclidean distance. This is illustrated in Fig. 2, which shows how the joint distribution captures unfairness, by depicting the percentage outreach in each group on each axis; thus, the diagonal represents a “fair” line, where the probability of reaching a particular outreach percentage is the same for both groups.
@@
where ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}z(x_{1},x_{2},y_{1},y_{2})}$ is the point indicated in green in Fig. 2 and $\norm{\cdot}$ is the standard Euclidean norm. Thus, the “fairness distance” between two distributions ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ can be readily quantified by $W_{c}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}})$. Since moving along the diagonal is free, we quantify the fairness of a given $\gamma$ as its “fairness distance" from the “ideal” distribution $\gamma^{\ast}=\delta_{(1,1)}$, which represents the case where all members of both groups receive the information. We can now formally introduce our proposed fairness metric.
@@
The mutual fairness from Definition 3.1 can be seen as a normalized expression of $W_{c}(\gamma,\gamma^{\ast})$ to contain its values in $[0,1]$. Indeed, its lowest value is 0 and it is achieved with $\gamma=\delta_{(0,1)}$, for which is $W_{c}(\gamma,\gamma^{\ast})=1$; its largest value is 1 and it is achieved with $\gamma=\gamma^{\ast}$, for which $W_{c}(\gamma^{\ast},\gamma^{\ast})=0$. Since $\gamma^{\ast}$ is a delta distribution, we can solve the transportation problem (1) in closed form to
@@ Discussion.
We note that while we considered two groups in the aforementioned definitions, our methodology readily extends the setting with $m$ groups. We present this extension in Appendix B. Second, since moving mass “diagonally” is free, any distribution $\gamma^{\ast}$ supported on the diagonal yields the same fairness metric. In practice, it is often not the case that all the network members receive the information and the best one could hope for is to project $\gamma$ onto the diagonal; since moving along the diagonal is free, the fairness cost is the same whether the ideal distribution is that projection or $\gamma^{\ast}$. Moreover, it is easy to see that the “fairness distance” is symmetric, namely $W_{c}(\gamma_{a},\gamma_{b})=W_{c}(\gamma_{b},\gamma_{a})$. Finally, our definition readily extends to any other distance function besides the standard Euclidean metric.
@@
which amounts to the cost of transporting the points $(0,1)$ and $(1,0)$, each with weight $1/4$, to the diagonal. Notably, in contrast to simply computing the expected outreach of each group, our fairness metric distinguishes the two outcomes. Similarly, we can easily compute the fairness metric: $\textsc{Fairness}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}})=1$ and $\textsc{Fairness}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}})=0.5$. In particular, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ achieves the highest fairness score. Indeed, its outcome will always be fair. Instead, $\textsc{Fairness}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}})$ achieves lower fairness score, capturing the fact that in 50% of the cases the outcome is perfectly fair while in the remaining 50% it is largely unfair.

### v4

@@ 1 Introduction
Social networks play a fundamental role in the spread of information, as in the context of commercial products endorsement [18], job vacancy advertisements [3], public health awareness [27], etc. Information, ideas, or new products can either go viral and potentially bring significant changes in a community or die out quickly. In this context, a fundamental algorithmic problem arises, known as Social Influence Maximization (SIM) [12, 13]. SIM studies how to strategically select a pre-specified small proportion of nodes in the social network – the early adopters or seeds – so that the outreach generated by a diffusion process that starts at these early adopters is maximized. Consider, for example, a product endorsement campaign: the early adopters are strategically selected users who receive the product first to promote it to their friends, who in turn may or may not adopt it. The optimal selection of early adopters is known to be an NP-hard problem [12]. Thus, many heuristic strategies have been proposed, based on iterative processes such as greedy algorithms or on network centrality measures. However, all these algorithms purely rely on the graph topology and are agnostic to users’ demographics, which raises significant fairness concerns, especially in contexts of health awareness campaigns, education, and job advertisements, where one wants to ensure an equitable spreading of information. Indeed, real-world social networks are populated by different social groups based on gender, age, race, geography, etc., with different group sizes or connectivity patterns. Ignoring these aspects and focusing only on the outreach maximization process usually leads to the early adopters being the most central nodes. Consequently, low-interconnected minorities are often neglected from the diffusion process, thus causing fundamental inequity in the information propagation and biases exacerbation [11, 25].
@@
The problem of SIM was first introduced in 2003 in Kempe et al. [12], where the problem of optimally selecting a (limited) set of early adopters was proved to be NP-hard. The study of SIM under fairness guarantees has a more recent history [6]. Several multiple group-level fairness metrics have been proposed over the years [7]. They fall under the notions of equity [23, 10, 11], equality [7], max-min fairness [8, 30], welfare [17], and diversity [25]: all of them quantify the fair distribution of influence across groups. In particular, Stoica et al. [23] propose a new SIM algorithm that operates under the constraint that, in expectation, the same percentage of users in each category is reached. Junaid et al. [10] optimize outreach under fairness and time constraints, by ensuring that the expected fraction of influenced nodes in each group is the same within a prescribed time deadline. Farnadi et al. [7] propose a unifying framework that encodes all different definitions of fairness in the SIM process as constraints in a linear program that optimizes outreach. Several other works [8, 30] adopt a max-min strategy. Specifically, in Fish et al. [8] fairness is ensured by maximizing the minimum probability of a group receiving the information through modifications of the greedy algorithm. Zhu et al. [30] ensure that the outreach contains a pre-specified proportion of each group in a population. Finally, Tsang et al. [25] optimize outreach under the constraint that no group should be better off by leaving the influence maximization process with their proportional allocation of resources done internally. All these definitions involve a marginal expected value of fairness in groups, without considering the correlations – or other higher-order moments – for the joint probability distribution of different groups adopting the information (see Farnadi et al. [7] for an overview). In contrast, our work introduces a novel formalism for taking into account the actual joint distribution of outreach among groups, thus considering all groups simultaneously, highlighting limitations of various fairness metrics and developing a new seed selection policy that strategically extracts and optimizes our proposed notion of fairness. To conclude, our work is inspired by a recent line of work that draws on optimal transport theory [28] for fairness guarantees [2, 4, 21, 29, 20, 24]. To our knowledge, this is the first work to develop novel metrics and seeding algorithms that leverage optimal transport for the SIM problem.
@@
Many models of diffusion processes in the SIM problem are inherently stochastic, meaning that who gets the information transmitted can vary greatly from one run to another. Consider, as an example, the case in which $50\%$ of realizations over a diffusion process, no one in group 1 receives the information and everyone in group 2 does, whereas in the other $50\%$ it is the opposite. This circumstance would be classified as fair in expectation, even though it is commonly not perceived as “fair”. We show how this phenomenon is common in real-world data and how our proposed framework can detect such undesired scenarios. This prompts us to study a novel fairness metric.
@@ 2 Preliminaries
For a network $G$ with two communities $(C_{i})_{i\in[2]}$ and a seedset $S$, we let $x_{i}$, $i\in[2]$, denote the fraction of nodes in each community in the outreach $\phi_{G}(S)$. The final configuration is the tuple $(x_{1},x_{2})$.
In many definitions in the literature, fairness is operationalized by measuring the expected value of the final configuration, where the expectation is taken over the diffusion process. In particular, the equity definition introduced by Stoica et al. [23], Junaid et al. [10] checks that the expected value of the proportions of each group reached in the outreach is the same for all groups. For a formal definition of equity and other fairness definitions in the literature, see Appendix A. We will show that relying solely on the expected value leads largely unfair outcomes to be classified as fair.
@@ 3 Mutual Fairness via Optimal Transport
Consider the SIM problem with nodes belonging to two groups, $C_{1}$ and $C_{2}$, each group having the outreach probability distribution $\mu_{i}=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{1},i\in\{1,2\}$, with $\delta_{k}$ representing the delta distribution at $k\in[0,1]$. That is, in $50\%$ of the cases all members in group $i$ receive the information (i.e., we get $x_{i}=1.0$) and in $50\%$ of the cases no one in group $i$ receives the information (i.e., we get $x_{i}=0.0$). It is therefore tempting to say that this setting is fair since $\mu_{1}$ and $\mu_{2}$ coincide and therefore share the same expected value. We argue that this information does not suffice to claim fairness. Indeed, consider the two following probability distributions over the final configurations:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}=0.5\cdot\delta_{(0,0)}+0.5\cdot\delta_{(1,1)},\qquad{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}=0.25\cdot\delta_{(0,0)}+0.25\cdot\delta_{(1,1)}+0.25\cdot\delta_{(0,1)}+0.25\cdot\delta_{(1,0)},$
with $\delta_{(i,j)}$, representing the delta distribution at $(i,j)\in[0,1]^{2}$. Interestingly, both ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ are “compatible” with $\mu_{1}$ and $\mu_{2}$: If we compute their marginals, we obtain $\mu_{1}$ and $\mu_{2}$. However, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ encode two fundamentally different final configurations. In ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$, the percentage of members of group 1 who get the information always coincides with the percentage of people of group 2. Conversely, in ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$, more outcomes are possible; in particular, there is a probability of $0.25+0.25=0.5$ that all members of one group receive the information and no member of the other group receives it (see Fig. 2). Thus, from a fairness perspective, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ encode very different outcomes. We therefore argue that a fairness metric should be expressed in terms of joint probability distribution $\gamma$, and not solely based on its marginals $\mu_{1}$ and $\mu_{2}$, as commonly done in the literature [23, 10].
@@ 3.1 A Fairness Metric Based on Optimal Transport
$W_{c}(\gamma_{a},\gamma_{b})=\min_{\pi\in\Pi(\gamma_{a},\gamma_{b})}\mathbb{E}_{(x_{1},x_{2}),(y_{1},y_{2})\sim\pi},[c((x_{1},x_{2}),(y_{1},y_{2}))],$ (1)
where $\Pi(\gamma_{a},\gamma_{b})$ is the set of probability distributions over $([0,1]\times[0,1])\times([0,1]\times[0,1])$ so that the first marginal is $\gamma_{a}$ and the second marginal is $\gamma_{b}$. Intuitively, the optimal transport problem quantifies the minimum transportation cost to morph $\gamma_{a}$ into $\gamma_{b}$ when transporting a unit of mass from $(x_{1},x_{2})$ to $(y_{1},y_{2})$ costs $c((x_{1},x_{2}),(y_{1},y_{2}))$. The optimization variable $\pi$ is called transportation plan and $\pi((x_{1},x_{2}),(y_{1},y_{2}))$ indicates the amount of mass at $(x_{1},x_{2})$ displaced to $(y_{1},y_{2})$. Thus, its first marginal has to be $\gamma_{a}(x_{1},x_{2})$ (that is, $(x_{1},x_{2})$ has to be transported to some $(y_{1},y_{2})$) and its second marginal must be $\gamma_{b}(y_{1},y_{2})$ (that is, the mass at $(y_{1},y_{2})$ has to arrive from some $(x_{1},x_{2})$). If the transportation cost $c$ is chosen to be a $p\geq 1$ power of a distance $d$, then $(W_{d^{p}}(\cdot,\cdot))^{1/p}$ is a distance on the space of probability distributions. When the probability distributions are discrete (or the space $[0,1]$ is discretized), the transportation problem (1) is a finite-dimensional linear program and can therefore be solved efficiently [16].
@@
To operationalize the optimal transport problem (1), we therefore need to define (i) a transportation cost and (ii) a reference distribution $\gamma^{\ast}$. To define the transportation cost, we start with the following two considerations. First, moving mass along the diagonal should have zero cost, as it does not affect fairness but only efficiency (the proportion of population reached in respective groups). Second, moving mass orthogonally towards the diagonal should come at a price, since the difference in group proportion outreach between groups 1 and 2 decreases. We quantify this price as the Euclidean distance. This is illustrated in Fig. 2, which shows how the joint distribution captures unfairness, by depicting the percentage outreach in each group on each axis; thus, the diagonal represents a “fair” line, where the probability of reaching a particular outreach percentage is the same for both groups.
@@
where ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}z(x_{1},x_{2},y_{1},y_{2})}$ is the point indicated in green in Fig. 2 and $\norm{\cdot}$ is the standard Euclidean norm. Thus, the “fairness distance” between two distributions ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}}$ can be readily quantified by $W_{c}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}})$. Since moving mass along the diagonal is free, we quantify the fairness of a given $\gamma$ as its “fairness distance" from the “ideal” distribution $\gamma^{\ast}=\delta_{(1,1)}$, which represents the case where all members of both groups receive the information. We can now formally introduce our proposed fairness metric.
@@
The mutual fairness from Definition 3.1 can be seen as a normalized expression of $W_{c}(\gamma,\gamma^{\ast})$ to contain its values in $[0,1]$. Indeed, its lowest value is 0 and it is achieved with $\gamma=\delta_{(0,1)}$, for which is $W_{c}(\gamma,\gamma^{\ast})=1$; its largest value is 1 and it is achieved with $\gamma=\gamma^{\ast}$, for which $W_{c}(\gamma^{\ast},\gamma^{\ast})=0$. Since $\gamma^{\ast}$ is a delta distribution, we can solve the optimal transport problem (1) in closed form to
@@ Discussion.
We note that, while we considered two groups in the aforementioned definition, our methodology readily extends the setting with $m$ groups. We present this extension in Appendix B. Second, since moving mass “diagonally” is free, any distribution $\gamma^{\ast}$ supported on the diagonal yields the same fairness metric. In practice, it is often not the case that all network members receive the information, and the best one could hope for is to project $\gamma$ onto the diagonal; since moving along the diagonal is free, the fairness cost is the same whether the ideal distribution is that projection or $\gamma^{\ast}$. Moreover, it is easy to see that the “fairness distance” is symmetric, namely $W_{c}(\gamma_{a},\gamma_{b})=W_{c}(\gamma_{b},\gamma_{a})$. Finally, our definition readily extends to any other distance function besides the standard Euclidean metric.
@@
which amounts to the cost of transporting the points $(0,1)$ and $(1,0)$, each with weight $1/4$, to the diagonal. Notably, in contrast to simply computing the expected outreach of each group, our fairness metric distinguishes the two outcomes. Similarly, we can easily compute the fairness metric: $\textsc{Fairness}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}})=1$ and $\textsc{Fairness}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}})=0.5$. In particular, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{a}}$ achieves the highest fairness score. Indeed, its outcome will always be fair. Instead, $\textsc{Fairness}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{b}})$ achieves a lower fairness score, capturing the fact that in 50% of the cases the outcome is perfectly fair, while in the remaining 50% it is largely unfair.